OpenStudy (anonymous):

evaluate (11^1/9)-^12

jhonyy9 (jhonyy9):

\[(11)^{-12/9} =(11)^{-4/3}= 1/(11)^{4/3}\]

jhonyy9 (jhonyy9):

\[1/11(11)^{1/3}\]

jhonyy9 (jhonyy9):

\[1/(11(11)^{1/3})\]

OpenStudy (anonymous):

sorry, i thought I deleted this question, its wrong

jhonyy9 (jhonyy9):

ok not is problem